If 484 J of energy is spent in increasing the speed of a wheel from 60rpm to 360 rpm the M.I is

Grade:11

1 Answers

HARDIK
6 years ago
STRATEGY:  SINCE 484J ENERGY IS SUPPLIED TO IT. THEN CHANGE IN ROTATIONAL KINETIC ENERGY IS EQUAL TO ENERGY SUPPLIED TO IT.
SOLUTION:  
(1/2)IWF2 – (1/2)IWi2 = 484 (WHERE WF AND WI  ARE INITIAL AND FINAL ANGULAR VELOCITIES)
WI = 6 RAD/S
WF = 37 RAD/S
HENCE PUTTING THE VALUES
(1/2)I(37*37)-(1/2)I(6*6)= 484
684.5I-18I = 484
HENCE MOMENT OF INERTIA= 0.72

REGARDS
HARDIK BAWEJA
PLEASE CHHECK CALCUTION,,,,,,,,